# “I use ChatGPT to humanize my words”: Affordances and Risks of ChatGPT to Autistic Users

### v1

@@ Abstract.
Large Language Model (LLM) chatbots like ChatGPT have emerged as cognitive scaffolding for autistic users, yet the tension between their utility and risk remains under-articulated. Through an inductive thematic analysis of 3,984 social media posts by self-identified autistic users, we apply the Technology Affordance framework to examine this duality. We found that while users leveraged ChatGPT to offload executive dysfunction, regulate emotions, translate neurotypical communication, and validate their autistic identity, these affordances coexist with significant risks: reinforcing delusional thinking, erasing authentic identity through automated masking, and triggering conflicts with the autistic sense of justice. This poster identifies these trade-offs in autistic users’ interactions with ChatGPT and concludes by outlining our future work on developing neuro-inclusive technologies that address these tensions through beneficial friction and bidirectional translation.
copyright: noneccs: Human-centered computing Empirical studies in HCI
@@ 1. INTRODUCTION
Large Language Model (LLM) chatbots like ChatGPT (OpenAI, n.d.) have quickly permeated everyday life. Beyond their functional utility, like information acquisition (Sun and Kalar, 2025) and writing support (Jelson et al., 2025), they are increasingly utilized for informal mental health support (Xu et al., 2024; Badawi et al., 2025), as users describe these chatbots as always-available listeners when professional care is geographically or financially inaccessible (Lawrence et al., 2024; Hua et al., 2025). For autistic adults [^1] We use identity-first language (e.g., “autistic person”) rather than person-first language. Prior research shows that identity-first language is preferred by the majority of the autistic community and research, as it reflects neurodiversity as a natural variation rather than a disorder (Bottema-Beutel et al., 2021; Kenny et al., 2016). We avoid the medicalized term autism spectrum disorder (ASD) unless explicitly citing clinical literature or describing our methods., LLM chatbots offer a personal and private supporting space. Emerging HCI work by Choi et al. (2024) highlights that users value ChatGPT for offering a “non-judgmental space” to tackle everyday chores, while Jang et al. (2024) found that autistic employees rely on GPT-4-based agents to navigate complex workplace dynamics. This work suggests that the appeal of ChatGPT lies in its ability to bridge the gap between the needs of individuals with autism and the demands of neurotypical societal interactions.
@@
However, such accessibility might obscure underlying tensions. Before the emergence of LLM chatbots, scholars had shown that autistic people’s engagement with AI technology is rarely purely beneficial. For example, on TikTok, autism content can help users reinterpret past experiences and engage in self-discovery (Alper et al., 2025), yet this simultaneously gives rise to “platformed diagnosis” a term coined by Alper and colleagues to describe how recommendation systems and diagnostic sense-making become mutually constitutive, raising bioethical concerns about algorithmic pseudo-diagnosis and the platform’s epistemic authority (Alper et al., 2025). Importantly, recent HCI scholarship has started to identify specific risks of LLM chatbots to autistic users, such as automating “masking [^2] Masking refers to the mentally demanding process of suppressing autistic traits to fit social norms, which relies heavily on executive functions like self-monitoring and inhibition (Hill, 2004).,” the suppression of authentic autistic traits, in hiring contexts (Hall et al., 2025) or creating intangible privacy vulnerabilities (Glazko et al., 2025), these findings often appear in isolation. The existing literature tends to bifurcate the user experience, treating the functional benefits (Choi et al., 2024; Jang et al., 2024) and the risks (Hall et al., 2025; Carik et al., 2025b) as relatively separate phenomena. We currently lack a more unified understanding of how the sociotechnical interaction between autistic users and LLM chatbots simultaneously generates these divergent outcomes.
@@
To capture this duality, we adopt the Technology Affordance framework. Originally defined by Gibson (Gibson, 2014) as action possibilities offered by the environment and “require both entities, but exist independent of the actor’s perceptions (Fox and McEwan, 2017).” Building on Gibson’s work, Don Norman introduced affordances to HCI and explicated both “real” and “perceived affordances (Norman, 1999)”. This study defined it as “the perceived and actual properties of the thing, primarily those fundamental properties that determine just how the thing could possibly be used” (Norman, 2013). We also argued for the importance of examining the “perceived” affordances, which acknowledge affordances as perceivable and allow a range of possibilities based on the user’s experience. The affordances approach acknowledges both the materiality of technology and human agency, enabling scholars to discuss technology without being accused of either technological determinism or social constructivism (Ellison et al., 2015). Therefore, it is well-suited to our study to analyze the relationship between autistic users and LLM chatbots, rather than focusing solely on their features. We aim to address the research question: What are the perceived affordances and associated risks of LLM chatbots for autistic users?
@@
To answer this research question, we conducted an inductive thematic analysis of a dataset of 3,984 social media posts (i.e., Reddit, X, and Tumblr) in which self-identified autistic users discussed their experiences with ChatGPT. We found that users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity. However, these affordances are inextricably linked to risks. The same algorithmic capabilities that provide accessibility simultaneously facilitate the reinforcement of delusional thinking, the displacement of authentic identity through automated masking, and ethical conflicts with the autistic sense of justice.
@@
This poster paves the way for rethinking neuro-inclusive design for LLM chatbots like ChatGPT, balancing cognitive support with autistic user agency. By adopting the Technology Affordance framework, we move beyond treating benefits and risks as isolated phenomena, revealing how certain capabilities of ChatGPT, as a representative LLM chatbot, bring them both to the autistic experience. We conclude by proposing directions for our future work, specifically the integration of beneficial friction and bidirectional translation tools, to mitigate the risks of dependency and identity erasure while preserving the essential cognitive scaffolding autistic users rely on.
@@ 2. METHODS
Data Collection. Understanding users’ experiences with digital technologies through social media data (e.g., Reddit) is a well-established method in HCI (e.g., (Ma and Kou, 2021; Zhang et al., 2025; Chen et al., 2025)). Using Brandwatch [^3] BrandWatch: https://www.brandwatch.com, we fetched posts from Reddit, X, and Tumblr between January 1, 2023, and September 30, 2025. We selected these platforms to maximize textual narrative depth compared to visual-first platforms like Instagram or YouTube, while this specific timeframe captures the rapid societal adoption following ChatGPT’s public release in late 2022 (Marr, 2023). Our search strategy focused on “ChatGPT,” chosen as the market leader to capture the most prevalent user experiences, though we acknowledge the growing ecosystem of other LLM chatbots, and further combined with autism-related keywords (e.g., “autism spectrum disorder,” “autistic,” “Asperger”) derived from the Diagnostic and Statistical Manual of Mental Disorders, Fifth Edition (DSM-5) (Diagnostic, 2013) and the validated Self-reported Mental Health Diagnoses (SMHD) dataset (Cohan et al., 2018). This targeted retrieval yielded a final dataset of 3,984 unique posts. Our university approved the study under the IRB exempt for secondary, anonymous online content.
@@
Data Processing. To ensure precision, we employed an LLM-assisted pipeline using GPT-4o-mini (OpenAI, 2025) to verify that posts explicitly discussed ChatGPT interactions in the context of autism. We implemented Chain-of-Thought (CoT) reasoning (Wei et al., 2023) to filter out unrelated noise. Two researchers independently coded a random sample of 162 posts to validate this pipeline, achieving strict consensus and high inter-rater reliability ($\alpha=1.00$ for ChatGPT relevance; $\alpha=0.91$ for autism relevance). The resulting 3,984 posts served as the corpus for our inductive thematic analysis.
@@
Data Analysis. We performed an inductive thematic analysis (Braun and Clarke, 2006, 2019) on the filtered corpus of 3,984 posts to examine how autistic users engage with ChatGPT. The first author familiarized themselves with the dataset before performing coding in Google Sheets. This yielded 239 initial codes of perceived affordances and 50 initial codes of perceived risks (see both codebooks in Appendix A). We coded data to capture both functional actions and user intent. For example, the statement, “I recognized part of the reason for my feelings was my Autistic drive for authenticity and justice,” was coded as “Conflict with sense of justice.” Following initial coding, the first author utilized Google Sheets to affinity diagram the codes, consolidating them into sub-themes and primary themes. All other co-authors met and discussed regularly with the first author to discuss the coding schemes, review quotations, and resolve disagreements. This single coder-led practice and team-based refinement align with thematic analysis practices in prior HCI work (e.g., (Jiang et al., 2021; Patel et al., 2019; Gauthier et al., 2022)).
@@ 3. FINDINGS
We found that autistic users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity. However, these perceived affordances coexist with risks of reinforcing delusional thinking, diminishing their identity, and conflicting with their sense of justice.
@@ 3.1.1. Offloading Executive Tasks to Get Unstuck
This theme captures the use of ChatGPT to bridge the gap between intent and action caused by executive dysfunction (Hill, 2004), characterized by difficulties in initiating tasks, maintaining focus, and organizing thoughts. Users turned to ChatGPT to generate momentum, as one user explained:
Ironically, ChatGPT is great for executive dysfunction because it lets you outsource your executive functions to a large language model. You don’t have to worry about the ‘starting’ or the structuring so much… you just dump the input, and it initiates the process for you.
Here, rather than struggling to structure ideas, a common barrier in executive dysfunction, this user “dumped” raw input directly into ChatGPT. ChatGPT substituted the missing executive planning step, converting chaotic thought processes into a coherent draft for refinement. Whether summarizing dense texts or structuring wandering thoughts, this externalization of executive function allowed autistic users to bridge the gap between intent and execution.
@@ 3.1.2. Relying on Non-Human Support for Emotional Regulation
We also found that users use ChatGPT as an on-demand outlet to manage acute anxiety without the unpredictable costs of human interaction (Mazefsky et al., 2013). Autistic users engaged ChatGPT to process intense emotions because it offered patience without the pressure to mask autistic traits (Hill, 2004). Unlike fatigued or reactive human support, ChatGPT offered a perceived safe container for processing meltdowns, confusing encounters, or repetitive inquiries where anxiety about straining relationships otherwise prevented help-seeking.
For example, users substituted human listeners with ChatGPT to avoid passive-aggressive responses. One user noted that “the only time I don’t feel some level of fear that I will accidentally annoy someone by asking too many questions,” identifying its lack of ego as a shield against social rejection. Utility lay not in ChatGPT’s ability to feel, but in its perceived neutrality, allowing users to regulate their emotional state through a source incapable of judgment.
3.1.3. Acting as a Translator for Neurodivergent-Neurotypical Communication
Additionally, autistic users described using ChatGPT to bridge the friction between neurodivergent communication styles and neurotypical social norms. Specifically, they employed ChatGPT to rewrite raw thoughts into professional norms, adding necessary social padding to blunt drafts to avoid perceived aggression. Conversely, they used ChatGPT to translate ambiguous messages by pasting confusing texts to detect sarcasm, subtext, or gaslighting. This allowed users to verify perceptions before responding, reducing anxiety over misinterpreting hidden social meanings.
For many, this translation proved essential for maintaining professional standing without the exhaustion of masking. One autistic user illustrated this by contrasting their internal monologue with ChatGPT’s output:
My autistic brain wants to say “Look, I contacted you about this 3 times in the past week… and all you did was say something vague about ‘later’ while my end users are furious”. Then ChatGPT spits out something like: “Hi W, Based on what I can see, my segment is functioning correctly given the parameters it’s receiving… I’ll need you to take the lead on resolving it.”
This quote illustrates that the user leveraged ChatGPT as a tonal translator, removing emotional urgency while preserving factual reality. Instead of struggling to manually encode frustration into acceptable office protocols, they used ChatGPT to convert a blunt demand into a neutral, boundary-setting statement.

### v2

@@ Abstract.
Large Language Model (LLM) chatbots like ChatGPT have emerged as cognitive scaffolding for autistic users, yet the tension between their utility and risk remains under-articulated. Through an inductive thematic analysis of 3,984 social media posts by self-identified autistic users, we apply a technology affordance lens to examine this duality. We found that while users leveraged ChatGPT to offload executive dysfunction, regulate emotions, translate neurotypical communication, and validate their autistic identity, these affordances coexist with risks to their well-being: reinforcing delusional thinking, erasing authentic identity through automated masking, and triggering conflicts with the autistic sense of justice. As part of our preliminary work, this poster identifies trade-offs in autistic users’ interactions with ChatGPT and concludes by outlining our future work on developing neuro-inclusive technologies that address these tensions through beneficial friction, bidirectional translation, and the delineation of emotional validation from reality.
journalyear: 2026copyright: ccconference: Interactive Health Conference; July 05–08, 2026; Porto, Portugalbooktitle: Interactive Health Conference (IH ’26), July 05–08, 2026, Porto, Portugaldoi: 10.1145/3786579.3804943isbn: 979-8-4007-2422-0/2026/07ccs: Human-centered computing Empirical studies in HCI
@@ 1. INTRODUCTION
Large Language Model (LLM) chatbots like ChatGPT (OpenAI, n.d.) have permeated everyday life. Beyond their functional utility (e.g., information acquisition (Sun and Kalar, 2025), writing support (Jelson et al., 2025)), they are increasingly utilized for informal mental health support (Xu et al., 2024; Badawi et al., 2025; Carik et al., 2025b), as users describe these chatbots as always-available listeners when professional care is inaccessible (Lawrence et al., 2024; Hua et al., 2025). For autistic adults [^1] We use identity-first language (e.g., “autistic person”) rather than person-first language. Prior research shows that identity-first language is preferred by the majority of the autistic community and research, as it reflects neurodiversity as a natural variation rather than a disorder (Bottema-Beutel et al., 2021; Kenny et al., 2016). We avoid the term autism spectrum disorder (ASD) unless citing clinical literature., LLM chatbots offer a personal and private supporting space. Relatedly, neurodivergent users describe LLMs as a supportive space for self-acceptance and emotional reflection (Carik et al., 2025b). Emerging work by Choi et al. (2024) highlights that users value ChatGPT as a “non-judgmental space” for everyday chores, while Jang et al. (2024) found that autistic employees rely on GPT-4-based agents to navigate workplace dynamics. Furthermore, recent HCI work leverages LLMs to decode conversational nuances for autistic users (Haroon and Dogar, 2024) and mediate balanced social interactions between neurodivergent and neurotypical individuals (Kong et al., 2025; Haroon et al., 2025; Carik et al., 2025b). This work suggests that the appeal of ChatGPT lies in its ability to bridge the gap between the needs of individuals with autism and the demands of neurotypical societal interactions, impacting their daily well-being.
@@
However, such accessibility might obscure underlying tensions. Before the emergence of LLM chatbots, scholars had shown that autistic people’s engagement with AI technology is rarely purely beneficial. On TikTok, the affordance of persistence (Treem and Leonardi, 2013; Hogan and Quan-Haase, 2010) regarding the AI-meidated autism content can help users reinterpret past experiences and engage in self-discovery (Alper et al., 2025), yet this simultaneously gives rise to “platformed diagnosis” describing how recommendation systems and diagnostic sense-making become mutually constitutive, raising bioethical concerns about algorithmic pseudo-diagnosis and the platform’s epistemic authority (Alper et al., 2025). Importantly, recent HCI scholarship has started to identify specific risks of LLM chatbots to autistic users, such as automating “masking [^2] Masking refers to the mentally demanding process of suppressing autistic traits to fit social norms, relying heavily on self-monitoring and inhibition (Hill, 2004).,” the suppression or misrepresentation of authentic autistic traits (Hall et al., 2025; Haroon and Dogar, 2024; Park et al., 2025), or creating intangible privacy vulnerabilities (Glazko et al., 2025). Building on this foundational research, which has naturally explored specific dimensions of this emerging phenomenon individually, we aim to extend the literature by examining both aspects together. Integrating the functional benefits (Choi et al., 2024; Jang et al., 2024) and the risks (Hall et al., 2025; Carik et al., 2025b) into a single analytical focus allows for a more unified understanding of how the sociotechnical interaction between autistic users and LLM chatbots simultaneously generates these divergent outcomes.
@@
To capture this duality, we adopt a technology affordance lens. Originally defined by Gibson (Gibson, 2014) as action possibilities offered by the environment and “require both entities, but exist independent of the actor’s perceptions (Fox and McEwan, 2017).” Building on this, Don Norman introduced affordances to HCI and explicated both “real” and “perceived affordances (Norman, 1999)”. Our study defined it as “the perceived and actual properties of the thing, primarily those fundamental properties that determine just how the thing could possibly be used” (Norman, 2013). We also argued for the importance of examining the “perceived” affordances, which acknowledge affordances as perceivable and allow a range of possibilities based on the user’s experience. Prior work in the non-LLM domain shows the value of this lens; for instance, examining how autistic individuals enact affordances on social media for self-presentation and affiliation (Koteyko et al., 2022), and how literal interpretations of platform affordances can lead to unintended negative consequences like social anxiety or exclusion (Page et al., 2022). The affordances approach acknowledges both the materiality of technology and human agency, enabling scholars to discuss technology without being accused of either technological determinism or social constructivism (Ellison et al., 2015). Therefore, it is well-suited to our study to analyze the relationship between autistic users and LLM chatbots, rather than focusing solely on their features. We ask: What are the perceived affordances and associated risks of LLM chatbots for autistic users?
@@
To answer this research question, we conducted an inductive thematic analysis of a dataset of 3,984 social media posts (i.e., Reddit, X, and Tumblr) that discuss autistic users’ experiences with ChatGPT. We found that self-identified autistic users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity. However, these affordances are inextricably linked to risks. The same algorithmic capabilities that provide accessibility simultaneously facilitate the reinforcement of delusional thinking, the displacement of authentic identity through automated masking, and ethical conflicts with the autistic sense of justice.
@@
This poster, as a part of our preliminary work on developing neuro-inclusive LLM chatbots, contributes to HCI and the health community by rethinking LLM chatbot design to balance cognitive support with autistic user agency, thereby safeguarding their psychological well-being. By adopting this technology affordance lens, we move beyond treating benefits and risks as isolated phenomena, revealing how ChatGPT simultaneously scaffolds and challenges the autistic experience. We conclude by proposing three directions for future work: integrating beneficial friction to prevent dependency, designing for interdependent, bidirectional translation to mitigate identity erasure, and delineating emotional validation from factual verification to avoid reinforcing delusional thinking.
@@ 2. METHODS
Data Collection. Understanding users’ experiences with digital technologies through social media data (e.g., Reddit) is a well-established method in HCI (e.g., (Ma and Kou, 2021; Zhang et al., 2025; Chen et al., 2025)). Using Brandwatch [^3] BrandWatch: https://www.brandwatch.com, we fetched posts from Reddit, X, and Tumblr between January 1, 2023, and September 30, 2025. We selected these platforms to maximize textual narrative depth compared to visual-first platforms like Instagram or YouTube, while this timeframe captures the rapid adoption following ChatGPT’s public release in late 2022 (Marr, 2023). Our search strategy required the simultaneous presence of the terms: “ChatGPT” AND an autism-related term (“autism” OR “autistic” OR “ASD” OR “Asperger’s”). We derived these keywords from the Diagnostic and Statistical Manual of Mental Disorders, Fifth Edition (DSM-5) (Diagnostic, 2013) and the validated Self-reported Mental Health Diagnoses (SMHD) dataset (Cohan et al., 2018). This initial Boolean retrieval yielded 6,013 unique posts for both GPT and autism. Our university approved the study under the IRB exempt for secondary online content.
@@
Data Processing. We employed an LLM-assisted pipeline using GPT-4o-mini (OpenAI, 2025) to verify that posts explicitly discussed ChatGPT interactions in the context of autism. We implemented Chain-of-Thought (CoT) reasoning (Wei et al., 2023) to filter out unrelated noise (e.g., AI news articles, third-party hypotheticals). Importantly, to ensure the dataset reflected first-person experiences, the CoT prompt instructed the model to retain only self-identified data. We acknowledge the methodological tension in using an LLM to filter a dataset about LLM usage. To mitigate potential algorithmic biases or self-preferential filtering from GPT-4o-mini, two researchers independently coded a random sample of 162 posts to validate this pipeline. Addressing the simultaneous presence of both topics (i.e., an AND condition for relevance), we achieved consensus and high inter-rater reliability ($\alpha=1.00$ for ChatGPT; $\alpha=0.91$ for autism relevance). This process removed 2,029 noisy posts, resulting in a relevant corpus of 3,984 posts that discussed both ChatGPT and autism for our inductive thematic analysis.
@@
Data Analysis. We performed an inductive thematic analysis (Braun and Clarke, 2006, 2019) on the filtered corpus of 3,984 posts to examine how autistic users engage with ChatGPT. The first author familiarized themselves with the dataset before performing coding in Google Sheets. During this initial coding phase, the researcher manually excluded posts discussing third-party experiences (e.g., “my husband is autistic”). We coded data directly related to our research question on perceived affordances and risks, ensuring the final thematic structure reflected only first-person, self-identified accounts. This yielded 239 unique initial codes for perceived affordances and 50 ones for perceived risks (see both codebooks in Appendix A). We coded data to capture both functional actions and user intent. For example, the statement, “I recognized part of the reason for my feelings was my Autistic drive for authenticity and justice,” was coded as “Conflict with sense of justice.” Following initial coding, the first author utilized Google Sheets to affinity diagram the codes, consolidating them into sub-themes and primary themes. All other co-authors met and discussed regularly with the first author to discuss the coding schemes, review quotations, and resolve disagreements. This single coder-led practice and team-based refinement align with thematic analysis practices in prior HCI work (e.g., (Jiang et al., 2021; Patel et al., 2019; Gauthier et al., 2022)).
@@ 3. FINDINGS
We found that autistic users leveraged ChatGPT to offload executive tasks, regulate emotions, translate neurodivergent-neurotypical communication, and validate their autistic identity. However, these affordances coexist with psychological risks of reinforcing delusional thinking, diminishing their autistic identity, and conflicting with their sense of justice.
@@ 3.1.1. Offloading Executive Tasks to Get Unstuck
This theme captures the use of ChatGPT to bridge the gap between intent and action caused by executive dysfunction (Hill, 2004), characterized by difficulties in initiating tasks, maintaining focus, and organizing thoughts. Users turned to ChatGPT to generate momentum, as one user explained: “Ironically, ChatGPT is great for executive dysfunction because it lets you outsource your executive functions to a large language model. You don’t have to worry about the ‘starting’ or the structuring so much… you just dump the input, and it initiates the process for you.” Here, rather than struggling to structure ideas, a common barrier in executive dysfunction, this user “dumped” raw input directly into ChatGPT. ChatGPT substituted the missing planning step; as another user noted, “I put the rambles of my autistic brain in there, and it makes it make sense.” Others relied on ChatGPT to overcome initiation barriers, with one sharing, “I was having trouble with task paralysis and threw everything I needed to do into ChatGPT and asked it to help me prioritize.” Whether summarizing dense texts or structuring wandering thoughts, this externalization allowed autistic users to bridge the gap between intent and execution.
@@ 3.1.2. Relying on Non-Human Support for Emotional Regulation
While prior work characterizes neurodivergent LLM use as a casual, “non-judgmental listener” (Carik et al., 2025b), our findings reveal a more important role, where it helped manage acute crises without the unpredictable costs of human interaction (Mazefsky et al., 2013). Autistic users engaged ChatGPT to process intense emotions in real-time; as one shared, “I just typed ’at public pool, overwhelmed. I’m autistic. Help’ and it HELPED.” It offered infinite patience without the pressure to mask (Hill, 2004), allowing users to communicate without fear of offense: “there’s also the refreshing way I don’t have to mask with an AI. It won’t get offended if I reply too matter-of-factly.” Unlike fatigued human support, ChatGPT provided a safe container for repetitive inquiries where anxiety about straining relationships otherwise prevented help-seeking. Users substituted human listeners to avoid social rejection and caregiver burnout, noting, “I reach to ChatGPT to help me work through difficult emotions. This also means that I don’t overburden the people I love.” Here, utility lay not in ChatGPT’s ability to feel, but in its perceived neutrality, allowing users to regulate their emotional state through a source incapable of judgment.
3.1.3. Acting as a Bi-directional Translator for Neurodivergent- Neurotypical Communication
While recent LLM systems aim to relieve autistic individuals from the one-sided burden of adapting to neurotypical norms (Haroon et al., 2025), autistic users still bear this communicative labor daily, using ChatGPT bi-directionally. To encode thoughts without perceived aggression, they added social padding to blunt drafts, with one noting, “I use ChatGPT to humanize my words in written conversations to sound less like an AI.” Conversely, they decoded ambiguous messages by pasting confusing texts to detect subtext. As one explained, “When I encounter manipulation, especially subtle gaslighting, I turn to ChatGPT to help me distinguish reality from distortion.” For many, this bi-directional translation proved essential for maintaining professional standing without the exhaustion of masking. One user contrasted their internal monologue with ChatGPT’s output: “My autistic brain wants to say “Look, I contacted you about this 3 times in the past week… and all you did was say something vague about ‘later’ while my end users are furious”. Then ChatGPT spits out something like: “Hi W, Based on what I can see, my segment is functioning correctly given the parameters it’s receiving… I’ll need you to take the lead on resolving it.” This case illustrates that the user leveraged ChatGPT as a tonal translator. Instead of struggling to manually encode frustration into acceptable office protocols, they used ChatGPT to convert a blunt demand into a neutral, boundary-setting statement.